# Reconciling Inflation with Hubble Anisotropies

### v1

@@ ABSTRACT
There have been persistent suggestions, based on several diverse data sets, that the cosmic expansion is not exactly isotropic. It is not easy to develop a coherent theoretical account of such a “Hubble anisotropy”, for, in standard General Relativity, intuition suggests that it contradicts the predictions of the very successful Inflationary hypothesis. We put this intuition on a firm basis, by proving that if we [a] make use of an Inflationary theory in which Inflation isotropises spatial geometry —  this, of course, includes the vast majority of such theories —  and if [b] we insist on assuming that spacetime has a strictly metric geometry (one in which the geometry is completely determined by a metric tensor), then indeed all aspects of the “Hubble field” must be isotropic. Conversely, should a Hubble anisotropy be confirmed, then either we must contrive to build anisotropy into Inflation from the outset, or we will have to accept that spacetime geometry is not strictly metric. We argue that allowing spacetime torsion to be non-zero would be by far the most natural way to accommodate such observations. Such theories make firm predictions, as for example that there should be a correlation between the degree of anisotropy at the end of Inflation and a certain specific component of the Hubble tension.
@@ 1. Inflation and the Hubble Field
However, Inflation is still called upon to explain (or at least render less mysterious) some much more basic observations. Among these is the strange fact that, while spacetime has an observably non-trivial geometry, it seems that, at cosmological scales, three-dimensional space does not. Much effort has been expended on searches for the curvature of three-dimensional space, to no avail thus far (see [4] for a clear discussion, and [5, 6] for the latest data on this).
We are entitled to ask why our spacetime has this extraordinary structure. And it is indeed extraordinary: the only way for the spatial sections to have an effectively Euclidean [^1] We use “Euclidean” to mean “devoid of non-trivial geometry”, avoiding the word “flat” because it has misleading connotations: we mean that the curvature is zero, but also that any other measure of geometric non-triviality in three-dimensional space, such as torsion, should vanish. intrinsic geometry is if the spacetime curvature, evaluated on them, resides entirely in their extrinsic geometry. But, for technical reasons to be briefly reviewed later (see the footnote in Section 3, below), this means that our spacetime is foliated by spatial slices in a very unusual way, one not possible for a generic spacetime. This calls for an explanation. (Note that this is not the traditional “Flatness Problem”: we are referring to the highly non-generic manner in which spacetime is foliated at the present time.)
Inflation explains this strange structure by, first, supposing that the spatial curvature in the pre-Inflationary Universe was bounded; this is motivated by the standard hopes about the avoidance of singularities in some future complete theory of the earliest times. It is then plausible that the enormous, de Sitter-like expansion characteristic of Inflation reduced the region of space we can observe to such a small subspace (see [7] for this) that it becomes indeed effectively Euclidean. That sets up just the right initial conditions for the part of spacetime we can observe to evolve to its present unusual state, with spatial sections dominated by extrinsic curvature.
@@
Now it should be noted that making this standard argument that “Inflation implies isotropy” reasonably rigorous is surprisingly difficult: see [8] for a good discussion, particularly of the problems associated with the quantum-mechanical aspects. And indeed, various aspects of this argument have, from time to time, been questioned. For example, it has sometimes been asked, in effect, whether Inflation really gives rise to enough expansion to “inflate away” the geometry of three-dimensional space: see for example [9] and particularly [10]. From a completely different viewpoint, it has been pointed out that the field responsible for Inflation, the “inflaton”, does not itself “inflate away”; if, therefore, the inflaton is inherently anisotropic —  if, for example, it is a vector field —  then we might not expect Inflation to predict statistical isotropy (see the discussion at the end of [8]).
However, these alternatives to the standard picture of Inflation, while interesting, seem rather contrived; in the case of “vector Inflation”, it is also difficult to arrange for the field to be stable. It seems to us that the peculiar spacetime structure we observe, with essentially all of the spacetime curvature (evaluated on spatial sections) concentrated in the extrinsic curvature of those sections, reflects the action of some profound physical process, and is no accident. We shall therefore, henceforth, cleave to the standard picture, as described above: we will take it that Inflation does indeed “inflate away” all non-trivial spatial geometry, so that the spatial sections are effectively exactly Euclidean, and therefore statistically isotropic. This approach also has the advantage of simplicity (and clarity).
All this has recently become topical and relevant, because statistical isotropy has been cast into doubt by analyses of several data sets: see for example [11, 12, 13, 14]. (See also Section 4.6 of [15] for a survey of these and other possible anomalies in the recent cosmic observations.) While this “statistical anisotropy” is by no means confirmed [16], recent observational advances promise to settle the question in the not very distant future. Since these new observations challenge the very foundations of the standard understanding of the early Universe, they certainly deserve attention.
@@
These remarkable new developments might well give us effective ways of studying statistical anisotropy, independently of the redshift field (which of course remains of great interest, and in which anisotropies may also exist, see for example [21] and especially the very clear presentation in [22]). (Thus the plural in the title of the present work.)
@@
To summarise: the fate of the Hubble field is connected to that of the intrinsic geometry, including its isotropisation, in rather subtle ways. We wish to argue from this that it is not obvious that Inflationary isotropisation of spatial sections necessarily entails an isotropic Hubble field.
Our objective here is to understand this precisely. We will find that, in all conventional Inflationary theories (meaning those in which anisotropy is not built in from the outset), and in all theories of gravitation using strictly metric spacetime geometry (meaning geometry which is completely fixed by a semi-Riemannian metric and nothing else), the extrinsic geometry does necessarily share the symmetries of the intrinsic geometry: in all such theories, the Hubble field (and the matter fields) must isotropise if Inflation occurs. This is why it is so difficult to reconcile statistical anisotropy with Inflation.
@@
First, when torsion is present, it is indeed possible for intrinsic isotropy to co-exist with anisotropy of the Hubble field: in other words, torsion supplies a natural way of reconciling Inflation with (some) observations of statistical anisotropy.
@@
Third, the extent of the anisotropy induced by torsion is measured by parameters which appear again in the generalised Friedmann equation. Thus for example axial torsion theories imply the existence of correlation between the Hubble anisotropy and a possible torsional component of the much-discussed “Hubble tension” [31, 32, 33]; this means that observations of statistical anisotropy might in principle allow (falsifiable) predictions as to the value of that component, and vice versa.
Finally, torsion can and often does disrupt the familiar simple relation between cosmic redshift and the scale factor, leading in some cases to a very complex and subtle form of anisotropy in the redshift field.
In summary, we argue that torsion offers the most natural path to reconciling Inflation with (some) possible observations of statistical anisotropy, while linking them to other cosmological anomalies, such as the Hubble tension.
Before we begin, we should point out that we are concerned here exclusively with what might be called “geometric” anisotropies. One should bear in mind, however, the possibility of anisotropies due to other causes, such as those due to bulk flows [34, 35] or to the growth of non-linear structures [36]. These might well be just as important as geometric anisotropies, and they may be superimposed on the effects we discuss. We exclude them here only to avoid confusion. (See however Section 8, below, where we will see that this separation may not ultimately be fully tenable.)
@@ 2. What Inflation Does and Does Not
To reiterate: the spatial sections of our Universe, those in which (apart from small peculiar velocities) the galaxies are distributed uniformly, appear (at present, and in the directly observable past) to be Euclidean, to a good approximation [4, 5, 6]. And yet, cosmic spacetime is by no means close to being Minkowskian. Spacetime has non-trivial geometry, but space does not. Some physical process has imposed this structure on spacetime. We assert that that process is what we call “inflating away”.
@@ 3. Geometry of the Hubble Field
for vectors W,Z,X,YW,\,Z,\,X,\,Y tangential to $\Sigma$, and where $h$ is the (0, 2)(0,\,2) version of $\hat{h}$. Note (see [37]) that this expression for the extrinsic curvature is true for any spacetime geometry endowed with a spacetime metric and a compatible linear connection: it is valid whether or not the geometry is what we are calling “strictly metric”. (That is, it is valid even if $h$ is not a symmetric bilinear form; but in that case the ordering of the vectors in equation (9) must be carefully preserved.)
@@ 4. Isotropisation of the Hubble Field in Strictly Metric Inflation
Now, in Inflation, isotropy is approached asymptotically [8]; that is, we are not interested in a situation where isotropy holds only at a single moment in time. Instead, we want the geometry around $\gamma(t)$ also to be approximately isotropic for at least some interval, of non-zero length, of parameter values along $\gamma$. This means that the Killing vectors associated with the isotropic mappings must commute with $\mathcal{L}_{\xi}$. Therefore equation (10) implies that each isotropic mapping of $g$ is also an isotropic mapping of $h$, that is, of $\mathfrak{H},$ since the two are equal in the strictly metric case.
@@
In the simplest cases one can easily confirm that $H$ is as expected. If we choose a (proper) time coordinate such that the unit normal is $\xi=\partial_{t},$ and set $g=a(t)^{2}\,f,$ where $a(t)$ is the scale factor and $f$ is a time-independent isotropic three-dimensional spatial metric tensor, then we have
@@
Inflation of a strictly metric spacetime, then, is a process whereby the Hubble tensor grows (from a very small or zero initial value) to a tensor which, by the time reheating begins, has a large magnitude and is as isotropic as the spatial metric.
There are two remarkable aspects of this argument. The first is that we have used the word “Inflation” in its broadest possible sense: as referring to an early phase of explosive expansion. The argument therefore applies to all of the many conventional (scalar) Inflationary models, and so the conclusion cannot be avoided by tinkering with the details (the precise nature of the inflaton potential or other features of its Lagrangian, and so on).
@@
We wish to argue that option [a] should be rejected: the successes of conventional Inflation are too numerous and various. Therefore option [b] would be necessary if statistical anisotropy were to be confirmed observationally. The question is whether it is sufficient: does a non-strictly metric spacetime geometry permit a statistical anisotropy of some kind, even if Inflation has isotropised the intrinsic spatial geometry?
We will now argue that it does, and that in fact it does so in an extremely natural manner. Whether it can do so in a manner consistent with (all of) the observations is of course another matter. In fact, we will see that option [b] does permit some kinds of anisotropy, but, at least in the concrete examples we will study, not all.
@@ 5. Isotropisation and Torsion: General Theory
Like $h$, $\mathfrak{H}$ is not, generically [^12] One can contrive a situation in which the antisymmetric part of $\beta$ exactly cancels $\varkappa$, so that $\mathfrak{H}$ becomes symmetric. Clearly this is a very special, artificial arrangement, and the reader can readily verify that it cannot happen in the cases of physical interest we will study below., a symmetric bilinear form when the torsion is not zero. We saw earlier (Section 3, above) that this means that a non-trivial rotational component, $Q$ in equation (6), must be present. In short, a non-trivial rotational component of $\mathfrak{H}$ is a signal that torsion is present. This is in fact the simplest way of understanding the geometric significance of torsion in cosmology.
@@
The same reasoning applies here, and so, if we accept this line of argument, then both terms on the right of equation (22) must (independently) vanish at that time. Both terms will however grow as Inflation proceeds, so, in particular, the extrinsic torsion $\varkappa$ will not be negligible by the time of reheating: as we will see in detail later, it must grow, to compensate for the loss of the intrinsic torsion.
@@
We have no reason to think so; $\varkappa$ and $\beta$ are not controlled by the intrinsic geometry. In fact it can be shown (Appendix 5 of [40]) that a real antisymmetric bilinear form like $\varkappa$ can be isotropic, in odd numbers of dimensions (3 in our case here), only if it is zero. Thus we see that $h$ can be isotropic only if the extrinsic torsion vanishes. In other words, the extrinsic geometry is automatically anisotropic if the extrinsic torsion does not vanish.
Similarly, we have no reason to think that $\beta$ will isotropise; it certainly does not do so in the concrete examples we will consider [^13] One can however ask what happens if we force a torsional geometry to be highly symmetric. This question has led to important classification theorems: see [52, 53, 54, 55]. Our point here, however, is just that Inflation alone does not impose such a high degree of symmetry on spacetime..
@@
Equation (24) illustrates the situation very clearly. In a purely metric spacetime, the isotropisation of the spatial metric isotropises the Hubble field. In the presence of torsion, however, there are two distinct possible sources of anisotropy. First, the spatial sections will, if $\varkappa\neq 0,$ embed into spacetime in an anisotropic way, specified by the detailed structure of $\varkappa$. Secondly, if $\beta\neq 0,$ anisotropies in the “bulk” (regions of spacetime remote from us) can give rise to anisotropic effects in signals we observe, depending on the properties of $\beta$.

### v2

@@ ABSTRACT
There have been persistent suggestions, based on several diverse data sets, that the cosmic expansion is not exactly isotropic. It is not easy to develop a coherent theoretical account of such a “Hubble anisotropy”, for, in standard General Relativity, intuition suggests that it contradicts the predictions of the very successful Inflationary hypothesis. We put this intuition on a firm basis, by proving that if we [a] make use of an Inflationary theory in which Inflation isotropises spatial geometry —  this, of course, includes the vast majority of such theories —  and if [b] we insist on assuming that spacetime has a strictly metric geometry (one in which the geometry is completely determined by a metric tensor), then indeed all aspects of the “Hubble field” must be isotropic. Conversely, should a Hubble anisotropy be confirmed, then either we must contrive to build anisotropy into Inflation (and into the geometry of space) from the outset, or we will have to accept that spacetime geometry is not strictly metric. We argue that the second option, implemented by allowing spacetime torsion to be non-zero, would be by far the most natural way to accommodate such observations. Such theories can reconcile non-isotropic matter distributions with a perfectly isotropic spatial geometry, and thus are able to reconcile Inflation with possibly observed anisotropies. They also allow us to reconcile the absence of anisotropy in one era (say, that of the CMB) with its presence in another.
@@ 1. Inflation and the Hubble Field
However, Inflation is still called upon to explain (or at least render less mysterious) some much more basic observations. Among these is the strange fact that, while spacetime has an observably non-trivial geometry, it seems that, at cosmological scales, three-dimensional space does not [^1] The evidence for Euclidean spatial geometry is interpreted as such in the context of General Relativity. In this work, we propose to modify GR (a little), so, strictly speaking, we should not assert that there is strong evidence for Euclidean spatial geometry in the modified theory. However, it is not plausible that such a modified theory would conspire to mislead us, by somehow changing the CMB angular power spectrum, the BAO, and so on, in just such a manner as to cancel the effects of an actual non-trivial spatial geometry.. Much effort has been expended on searches for the curvature of three-dimensional space, to no avail thus far; it is important to note that this result has now been confirmed by many different lines of investigation [4, 5], with unprecedented precision (see [6, 7] for clear general discussions and [8, 9, 10] for the latest data on this).
We are entitled to ask why our spacetime has this extraordinary structure. And it is indeed extraordinary: the only way for the spatial sections to have an effectively Euclidean [^2] We use “Euclidean” to mean “devoid of non-trivial geometry”, (usually) avoiding the word “flat” because it has misleading connotations: we mean that the curvature is zero, but also that any other measure of geometric non-triviality in three-dimensional space, such as intrinsic torsion, should vanish. intrinsic geometry is if the spacetime curvature, evaluated on them, resides entirely in their extrinsic geometry. But, for technical reasons to be briefly reviewed later (see the footnote in Section 3, below), this means that our spacetime is foliated by spatial slices in a very unusual way, one not possible for a generic spacetime. This calls for an explanation. (Note that this is not the traditional “Flatness Problem”: we are referring to the highly non-generic manner in which spacetime is foliated at the present time.)
Inflation explains this strange structure by, first, supposing that the spatial curvature in the pre-Inflationary Universe was bounded; this is motivated by the standard hopes about the avoidance of singularities in some future complete theory of the earliest times. It is then plausible that the enormous, de Sitter-like expansion characteristic of Inflation reduced the region of space we can observe to such a small subspace (see [11] for this) that it becomes indeed effectively Euclidean. That sets up just the right initial conditions for the part of spacetime we can observe to evolve to its present state, with its unusual foliation.
@@
Now it should be noted that making this standard argument that “Inflation implies isotropy” reasonably rigorous is surprisingly difficult: see [12] for a good discussion, particularly of the problems associated with the quantum-mechanical aspects. And indeed, various aspects of this argument have, from time to time, been questioned. For example, it has sometimes been asked, in effect, whether Inflation really gives rise to enough expansion to “inflate away” the geometry of three-dimensional space: see for example [13] and particularly [14]. From a completely different viewpoint, it has been pointed out that the field responsible for Inflation, the “inflaton”, does not itself (necessarily) “inflate away”; if, therefore, the inflaton is inherently anisotropic —  if, for example, it is a vector field [15, 16] —  then we might not expect Inflation to predict statistical isotropy. (See the discussion at the end of [12]; see also [17] for some interesting recent developments in this field.)
This concept of “vector Inflation” builds anisotropy into Inflation from the outset, and arranges for that anisotropy to survive the extreme inflationary expansion. But it is a somewhat radical departure from Inflation as usually conceived, and, more importantly, it means that we have to abandon the claim, discussed above, that Inflation explains the fact that our Universe appears to be foliated by Euclidean spacelike hypersurfaces.
It seems to us that the peculiar spacetime structure we observe, with essentially all of the spacetime curvature (evaluated on spatial sections) concentrated in the extrinsic curvature of those sections, reflects the action of some profound physical process, and is no accident.
We shall therefore, henceforth, cleave to the standard picture, as described above: we will take it that Inflation does indeed “inflate away” all non-trivial spatial geometry, not approximately but essentially exactly, so that the spatial sections have an intrinsic geometry which is effectively exactly Euclidean, and therefore statistically isotropic. Our claim will be that, strange as it may seem, this picture can be reconciled with observations of cosmic anisotropy.
All this has recently become topical and relevant, because statistical isotropy has in fact been cast into doubt by analyses of several data sets: see for example [18, 19, 20, 21]. (See also Section 4.6 of [22] for a survey of these and other possible anomalies in the recent cosmic observations.) While this “statistical anisotropy” is by no means confirmed [23], it is difficult to believe that the agreement of all these diverse data is a “fluke” [24]. In any case, recent observational advances promise to settle the question in the not very distant future. Since these new observations challenge the very foundations of the standard understanding of the early Universe, they certainly deserve attention.
@@
These remarkable new developments might well give us effective ways of studying statistical anisotropy, independently of the redshift field (which of course remains of great interest, and in which anisotropies may also exist, see for example [29, 30, 31, 32, 33, 34, 35] and especially the very clear presentation in [36]). Thus the plural in the title of the present work.
@@
Our objective here is to understand this precisely. We will find that, in all conventional Inflationary theories (meaning those in which anisotropy is not built in from the outset), and in all theories of gravitation using strictly metric spacetime geometry (meaning geometry which is completely fixed by a semi-Riemannian metric and nothing else), the extrinsic geometry does necessarily share the symmetries of the intrinsic geometry: in all such theories, the Hubble field (and the matter fields) must isotropise if Inflation occurs. This is why it is so difficult to reconcile statistical anisotropy with (conventional) Inflation.
@@
First, when torsion is present, it is indeed possible for a perfectly isotropic intrinsic spatial geometry to co-exist with anisotropy of the matter distribution and therefore of the Hubble field: in other words, torsion supplies a natural way of reconciling Inflation with (some) observations of statistical anisotropy. (The essential technical point here is that the symmetries of the matter field are correlated, via the constraint equations, with the extrinsic, not the intrinsic, geometry of the spatial sections.)
@@
Third, the extent of the anisotropy induced by torsion is measured by parameters which appear again in the generalised Friedmann equation. Thus for example axial torsion theories imply the existence of correlation between the Hubble anisotropy and a possible torsional component of the much-discussed “Hubble tension” [45, 46, 47, 48]; this means that observations of statistical anisotropy might in principle allow (falsifiable) predictions as to the value of that component, and vice versa.
Finally, we will see that, in torsional theories, the level of anisotropy at one cosmic epoch (for example, as seen in the CMB) may not dictate the level observable at other times, potentially removing any conflict between data sets corresponding to differing eras.
Before we begin, we should point out that we are concerned here exclusively with what might be called “geometric” anisotropies. One should bear in mind, however, the possibility of anisotropies in the observations due to other causes, such as those due to bulk flows [49, 50] or to the growth of non-linear structures [51]. These might well be just as important as geometric anisotropies, and they may be superimposed on the effects we discuss. We exclude them here only to avoid confusion.
@@ 2. What Inflation Does and Does Not
To reiterate: the spatial sections of our Universe, those in which (apart from small peculiar velocities) the galaxies are distributed uniformly, appear (at present, and in the directly observable past) to be Euclidean, to a good approximation [6, 7, 8, 9, 10]. And yet, cosmic spacetime is by no means close to being Minkowskian. Spacetime has non-trivial geometry, but space does not. Some physical process has imposed this structure on spacetime. We assert that that process is what we call “inflating away”.
@@ 3. Geometry of the Hubble Field
for vectors W,Z,X,YW,\,Z,\,X,\,Y tangential to $\Sigma$, and where $h$ is the (0, 2)(0,\,2) version of $\hat{h}$. Note (see [52]) that this expression for the extrinsic curvature (and also the Gauss formula relating it to the full spacetime curvature) is true for any spacetime geometry endowed with a spacetime metric and a compatible linear connection: it is valid whether or not the geometry is what we are calling “strictly metric”. (That is, it is valid even if $h$ is not a symmetric bilinear form; but in that case the ordering of the vectors in equation (9) must be carefully preserved.)
@@ 4. Isotropisation of the Hubble Field in Strictly Metric Inflation
Now, in Inflation, isotropy is approached asymptotically [12]; that is, we are not interested in a situation where isotropy holds only at a single moment in time. Instead, we want the geometry around $\gamma(t)$ also to be approximately isotropic for at least some interval, of non-zero length, of parameter values along $\gamma$. This means that the Lie derivatives with respect to the Killing vectors associated with the isotropic mappings must commute with $\mathcal{L}_{\xi}$. Therefore equation (10) implies that each isotropic mapping of $g$ is also an isotropic mapping of $h$, that is, of $\mathfrak{H},$ since the two are equal in the strictly metric case.
@@
In the simplest cases one can easily confirm that $H$ is as expected. If we use Gaussian normal coordinates with a proper time coordinate $t$ (so that $\xi=\partial_{t}$ when both are evaluated on $\Sigma$) and set $g=a(t)^{2}\,f,$ where $a(t)$ is the scale factor and $f$ is a time-independent isotropic three-dimensional spatial metric tensor, then we have
@@
Inflation in the case of a strictly metric spacetime, then, is a process whereby the Hubble tensor grows (from a very small or zero initial value) to a tensor which, by the time reheating begins, has a large magnitude and is as isotropic as the spatial metric.
There are two remarkable aspects of this argument. The first is that we have used the word “Inflation” in its broadest possible sense: as referring to an early phase of extreme expansion. The argument therefore applies to all of the many conventional Inflationary models, and so the conclusion cannot be avoided by tinkering with the details (the precise nature of the inflaton potential or other features of its Lagrangian, and so on).
@@
We wish to argue that pursuing option [a] should be attempted only if all else fails. If we accept this, then option [b] would be necessary if statistical anisotropy were to be confirmed observationally. The question is whether it is sufficient: does a non-strictly metric spacetime geometry permit a statistical anisotropy of some kind, even if Inflation has isotropised the intrinsic spatial geometry?
We will now argue that it does, and that in fact it does so in an extremely natural manner. Whether it can do so in a manner consistent with (all of) the observations is of course another matter. In fact, we will see that option [b] does permit some kinds of anisotropy, but not all.
@@ 5. Isotropisation and Torsion: General Theory
Like $h$, $\mathfrak{H}$ is not, generically [^13] One can however force it to be symmetric by imposing isotropy “by hand”: see below., a symmetric bilinear form when the torsion is not zero. We saw earlier (Section 3, above) that this means that a non-trivial rotational component, $Q$ in equation (6), must be present. In short, a non-trivial rotational component of $\mathfrak{H}$ is a signal that torsion is present. This is in fact the simplest way of understanding the geometric significance of torsion in cosmology.
@@
The same reasoning applies here, and so, if we accept this line of argument, then both terms on the right of equation (22) must (independently) vanish at the beginning. Both terms will however grow as Inflation proceeds, so, in particular, the extrinsic torsion $\varkappa$ will not be negligible by the time of reheating: as we will see in detail later, it must grow, to compensate for the loss of the intrinsic torsion.
@@
We have no reason to think so; $\varkappa$ and $\beta$ are not controlled by the intrinsic geometry. In fact it can be shown (Appendix 5 of [55]) that a real antisymmetric bilinear form like $\varkappa$ can be isotropic, in odd numbers of dimensions (3 in our case here), only if it is zero. Thus we see that $h$ can be isotropic only if the extrinsic torsion vanishes. In other words, the extrinsic geometry is automatically anisotropic if the extrinsic torsion does not vanish. (A similar argument implies that the intrinsic torsion has to vanish if it is to be isotropic.)
Similarly, we have no reason to think that $\beta$ will isotropise; it certainly does not do so in the concrete examples we will consider. If we force it to be isotropic, then, as we have been discussing, its antisymmetric part has to vanish, and Schur’s lemma implies that its symmetric part is a time-dependent multiple, let us call it $4\phi(t)$, of $g$. If both the extrinsic and the “bulk” torsion are forced to be isotropic, then, both $h$ and $\mathfrak{H}$ are symmetric bilinear forms. (Note that the “bulk” torsion one-form $b$, given by equation (18), must vanish in the isotropic case because it is metrically dual to a vector field.)
This special case of “isotropic torsion”, in which all forms of torsion vanish except for the “bulk” torsion bilinear form $\beta,$ given here by $\beta=4\phi g,$ is the case usually assumed in discussions of torsional cosmology: see [67]. The question as to what happens if we force spacetime torsion to share the symmetries of the metric has been studied in a more general and complete manner, leading to important classification theorems: see [68, 69, 70, 71]. Our point here, however, is just that Inflation alone does not impose such high degrees of symmetry on spacetime torsion.
@@
Equation (24) illustrates the situation very clearly. In a purely metric spacetime, the isotropisation of the spatial metric isotropises the Hubble field. In the presence of torsion, however, there are two distinct possible sources of anisotropy. First, the spatial sections will, if $\varkappa\neq 0,$ embed into spacetime in an anisotropic way, specified by the detailed structure of $\varkappa$. Secondly, if $\beta\neq 0,$ anisotropies in the “bulk” (regions of spacetime remote from us) can and normally will give rise to anisotropies, which (as we will see later) can appear in signals we observe, arriving from afar.
@@
In the exceptional case of “isotropic torsion”, this equation takes the form $\mathfrak{H}=\left(H\,+\,2\phi\right)g$. (The coefficient of $\phi$ was chosen so as to ensure agreement with [67].) Thus “isotropic torsion” makes its presence felt through a modification of the Hubble parameter. We will see later that a similar effect arises in the anisotropic case.